# Amps: ASR with Multimodal Paraphrase Supervision

### v1

@@ Abstract
Amruta ${}^{*}\quad$ Abhishek ${}^{*}\quad$ Sameep Chattopadhyay    Preethi Jyothi Indian Institute of Technology Bombay, Mumbai, India {amrutaparulekar.iitb,abhishekumgupta,sameep.ch.2002}@gmail.com, pjyothi@cse.iitb.ac.in
@@ 1 Introduction
Amps leverages the multimodal nature of SeamlessM4T by introducing a paraphrasing objective jointly with ASR. Along with using spontaneous speech and its corresponding transcription to train the speech-to-text pathway in SeamlessM4T, Amps also uses paraphrases of the reference transcriptions as additional supervision to train the text-to-text pathway. We selectively employ paraphrase-based augmentation during training when the ASR loss is high (as determined by a predetermined threshold); high ASR loss is typically triggered by noise or poorly enunciated words in spontaneous speech. This selective intervention offers the model an alternate path of opting for semantically close words and phrases when the audio is not very clear. It is important that the paraphrases should not significantly differ in word order from the original transcripts, thus enabling the model to easily align representations of speech, text, and its paraphrase.
@@ 4 Experimental Setup
The IndicVoices dataset Javed et al. (2024b) is a large collection of natural speech (74% extempore, 17% conversational and 9% read) in 22 Indic languages. Among the languages we chose, Marathi, Kannada, and Malayalam are classified as low-resource by SeamlessM4T Communication et al. (2023), while Hindi is medium-resource. IndicVoices is the only multilingual open-source Indian speech corpus containing spontaneous speech and amongst the very few sources published after SeamlessM4T’s release.***To ensure that there was no data leakage between the SeamlessM4T training data and the evaluation sets. We also performed experiments on Nyanja (a low-resource language from Zambia) from the Zambezi-Voice dataset Sikasote et al. (2023).
@@ 5 Experiments and Results
Table 1 shows the main results for all the 50-hour Indian-language experiments. Ampsτ consistently performs best compared to ASR, and the WER reductions are statistically significant (at $p<0.05$ using the mapsswe test). Apart from the overall scores in All Data, we sorted the transcriptions in descending order of WER using pure ASR and averaged metrics were calculated for both pure ASR and Ampsτ for the first 100 (hardest) sentences. Improvements from ASR to Ampsτ for these hardest 100 predictions are labeled $\Delta$Hard in Table 1. We see that $\Delta$Hard consistently exceeds $\Delta$All, indicating that the most improvement is observed in cases where pure ASR performs poorly. This supports the thresholding approach that triggers the paraphrase loss only when pure ASR predictions fall below a threshold.
@@ 5.1 Comparing Paraphrase Techniques
Table 2 shows results from training on 5 hrs of read/mixed Hindi speech and different paraphrasing techniques with mixed speech. Unsurprisingly, training on mixed speech yields significantly lower WERs compared to training on read speech. The highest performance gains were obtained using LLM paraphrasing for Hindi, suggesting that the LLM is a good option for medium-resource languages like Hindi. LLM outputs are subpar for low-resource languages like Kannada, and hence not an option. Comprehensive results comparing the paraphrase techniques for other languages are given in Appendix F and G.
@@ 5.3 Amps for Nyanja
Table 4 shows overall results***WER and METEOR reported, BERTScore does not support Nyanja. on Nyanja with 5 hours of training data and round-trip translated paraphrases. Again, Ampsτ performs the best, showing that Amps could be applied to diverse languages across language families.
@@ 5.4 Conclusion
This work introduces a novel paraphrase-based supervision technique Amps to improve the ASR performance of spontaneous speech in multimodal models. This auxiliary supervision makes the model more robust and helps the model generalize better, especially in utterances with large ASR errors. We show significant ASR improvements on multiple and diverse languages and further validate these improvements via a thorough human evaluation. Future work will investigate how techniques like Amps could be used to improve ASR for atypical speech.

### v2

@@ Abstract
Abhishek ${}^{*}\quad$ Amruta ${}^{*}\quad$ Sameep Chattopadhyay    Preethi Jyothi Indian Institute of Technology Bombay, Mumbai, India {abhishekumgupta,amrutaparulekar.iitb,sameep.ch.2002}@gmail.com, pjyothi@cse.iitb.ac.in
@@ 1 Introduction
Amps  [^1] Code for AMPS is available at https://github.com/csalt-research/amps-asr. leverages the multimodal nature of SeamlessM4T by introducing a paraphrasing objective jointly with ASR. Along with using spontaneous speech and its corresponding transcription to train the speech-to-text pathway in SeamlessM4T, Amps also uses paraphrases of the reference transcriptions as additional supervision to train the text-to-text pathway. We selectively employ paraphrase-based augmentation during training when the ASR loss is high (as determined by a predetermined threshold); high ASR loss is typically triggered by noise or poorly enunciated words in spontaneous speech. This selective intervention offers the model an alternate path of opting for semantically close words and phrases when the audio is not very clear. It is important that the paraphrases should not significantly differ in word order from the original transcripts, thus enabling the model to easily align representations of speech, text, and its paraphrase.
@@ 4 Experimental Setup
The IndicVoices dataset Javed et al. (2024b) is a large collection of natural speech (74% extempore, 17% conversational and 9% read) in 22 Indic languages. Among the languages we chose, Marathi, Kannada, and Malayalam are classified as low-resource by SeamlessM4T Communication et al. (2023), while Hindi is medium-resource. IndicVoices is the only multilingual open-source Indian speech corpus containing spontaneous speech and amongst the very few sources published after SeamlessM4T’s release. [^2] This dataset was chosen also to ensure that there was no data leakage between the SeamlessM4T training data and the evaluation sets. We also performed experiments on Nyanja (a low-resource language from Zambia) from the Zambezi-Voice dataset Sikasote et al. (2023).
@@ 5 Experiments and Results
Table 1 shows the main results for all the 50-hour Indian-language experiments. Ampsτ consistently performs best compared to ASR, and the WER reductions are statistically significant (at $p<0.05$ using the mapsswe test). [^3] We also trained a variant where instances with a ASR loss were downweighted and instances with a high ASR loss were upweighted, thus forcing the model to focus more on the latter. This performed comparably to our baseline ASR model. Apart from the overall scores in All Data, we sorted the transcriptions in descending order of WER using pure ASR and averaged metrics were calculated for both pure ASR and Ampsτ for the first 100 (hardest) sentences. Improvements from ASR to Ampsτ for these hardest 100 predictions are labeled $\Delta$Hard in Table 1. We see that $\Delta$Hard consistently exceeds $\Delta$All, indicating that the most improvement is observed in cases where pure ASR performs poorly. This supports the thresholding approach that triggers the paraphrase loss only when pure ASR predictions fall below a threshold. From our manual inspection of Hindi samples in the hardest-100 subset, we observe examples where pure ASR tends to produce acoustically similar but incorrect words, while Ampsτ correctly identifies the words. For example, pure ASR misrecognized “hua" (meaning ’is’) as “ugwa" (meaning ’grows’) in a Hindi example; Ampsτ gets this example right.
@@ 5.1 Comparing Paraphrase Techniques
Table 2 shows results from training on 5 hrs of read/mixed Hindi speech and different paraphrasing techniques with mixed speech. Here, by mixed speech, we refer to a mixture of both read and conversational speech. Unsurprisingly, training on mixed speech yields significantly lower WERs compared to training on read speech. The highest performance gains were obtained using LLM paraphrasing for Hindi, suggesting that the LLM is a good option for medium-resource languages like Hindi. LLM outputs are subpar for low-resource languages like Kannada, and hence are not an option. Comprehensive results comparing the paraphrase techniques for other languages are given in Appendix F and G.
@@ 5.3 Amps for Nyanja
Table 4 shows overall results [^4] Only WER and METEOR are reported. BERTScore does not support Nyanja. on Nyanja with 5 hours of training data and round-trip translated paraphrases. Again, Ampsτ performs the best, showing that Amps could be applied to diverse languages across language families.
@@ 5.4 Conclusion
This work introduces a novel paraphrase-based supervision technique Amps to improve the ASR performance of spontaneous speech in multimodal models. This auxiliary supervision makes the model more robust and helps the model generalize better, especially in utterances with large ASR errors. We show significant ASR improvements on multiple and diverse languages and further validate these improvements via a thorough human evaluation.
The broader idea of using textual supervision, as we did with paraphrases, to improve speech understanding is an interesting avenue to explore further. Future work will investigate how techniques like Amps could be used to improve ASR for atypical speech. Also, we used a predefined threshold on the ASR loss to trigger the paraphrase objective; this could be made a learnable quantity.
6 Acknowledgements
The authors thank the anonymous reviewers for their constructive feedback that improved the quality of the draft. The last author gratefully acknowledges support from the Amazon IITB AI ML Initiative.